OpenStudy (stormoshawty):

Simplify the equation. 3^-2 / n

OpenStudy (anonymous):

since the 3^-2 it is equal to 1/9 so the answer should be 1/(9n)

OpenStudy (anonymous):

3^-2/n 1/3^2(n) 1/9n

OpenStudy (shamil98):

Negative exponents rule: x^(-n) = 1/x^n you have 3^-2 / n so 1/9 / n 1/9 * 1/n = 1/9n

OpenStudy (stormoshawty):

Thank you Shamil.

OpenStudy (shamil98):

np :)